OpenStudy (djbreezy):

Which is an equation in point-slope form for the given point and slope? Point: (-2,4) ; Slope: 5 A. y+4=5x-10 B. y+4=5(x+2) C. y-4=5(x+2) D.y-2=5(x-2)

OpenStudy (anonymous):

(y−y1)=m(x−x1) y1 = 4 m = 5 x1 = -2

OpenStudy (welshfella):

point slope form is y - y1 = m(x - x1) - where m = slope and (x1,y1) is the point on the line

OpenStudy (welshfella):

so plug in the values that french has written and simplify the right side

OpenStudy (welshfella):

- well - no need to simplify

OpenStudy (anonymous):

once you plug in those values you should get y-4=5(x+2)